# Motility-Driven Viscoelastic Control of Tissue Morphology in Presomitic Mesoderm

### v1

@@ IV Cell motility controls tissue viscoelasticity
Figure 2: Tissue rheology under standard mechanical protocols. (a) Stress relaxation response for tissues with different motility levels. The decay of shear stress $\sigma_{xy}(t)$ is shown following a step shear deformation. Higher motility leads to faster stress relaxation and lower residual stress. Inset: Long-time behavior of $\sigma_{xy}$ reveals a nonzero plateau, indicating residual stress retention. (b) Relaxation timescale $\tau_{\mathrm{s}}$ decreases with increasing motility, indicating enhanced fluidization. (c) Residual stress as a function of motility. More motile tissues retain less stress over time. (d) Frequency-dependent storage modulus $G^{\prime}$ and loss modulus $G^{\prime\prime}$. At low frequencies, viscous behavior dominates ($G^{\prime\prime}>G^{\prime}$); at higher frequencies, the tissue responds more elastically. (e) The loss tangent $\tan{\delta}=G^{\prime\prime}/G^{\prime}$ as a function of driving frequency $\omega_{0}$, illustrating the transition from viscous to elastic dominance. (f) Intrinsic timescale of the tissue($\tau_{v}$), defined by the crossover frequency where $G^{\prime}=G^{\prime\prime}$, shifts to lower values with increasing motility, reflecting faster stress relaxation dynamics in more active tissues.
@@
To further dissect tissue rheology, we subjected the tissue to controlled shear deformations under varying levels of cellular motility. This approach allowed us to quantify fundamental rheological properties such as stress relaxation dynamics and the frequency-dependent storage and loss moduli, thereby establishing a direct link between local active fluctuations and the emergent viscoelastic behavior at the tissue scale.
@@
IV.1 Stress Relaxation Test Reveals Motility Driven Residual Stress Retention
@@
The anterior–posterior variation in tissue viscosity suggests that different regions of the PSM may respond differently to mechanical loading. To investigate how this rheological heterogeneity affects stress dissipation, we performed stress relaxation tests on tissues with varying motility levels. In each simulation, we applied a step shear by deforming the tissue affinely and fixing the outermost cell layers to maintain the applied strain. We then tracked the temporal evolution of shear stress in the bulk region, away from the boundary. By comparing tissues with low, medium, and high motility representing anterior, middle, and posterior PSM, we examined how active cellular dynamics influence the tissue’s ability to dissipate stress.
@@
In all cases, we observed a characteristic biphasic relaxation behavior. At short to intermediate timescales, stress rapidly decayed (Fig. 2(a)). The slope of this early time relaxation phase was found to depend strongly on the cell motility: tissues with higher motility exhibited a steeper decay, indicating faster stress dissipation(Fig. 2(b)).
@@
At longer timescales, the stress curves plateaued, revealing a non-zero residual stress in the tissue [Fig. 2(a), Inset]. This behavior indicates solid-like features where cells cannot fully rearrange to relax all internal stresses. The magnitude of this residual stress was highest in the low-motility (anterior) regime and dropped as we went towards higher motility zones (posterior) [Fig. 2(c)]. This motility-dependent decline in residual stress qualitatively agrees with experimental observations in embryonic PSM where the posterior part shows lower residual stress than the anterior PSM (e.g. Kim et al. (2021); Mongera et al. (2018)).
@@
These stress relaxation test results suggest that motility not only governs the rate of stress dissipation but also tunes the mechanical state of the tissue from a more elastic, stress-retaining regime to a more fluidized, actively relaxing state. These findings help explain how developing tissues adjust their mechanical behavior during morphogenesis by tuning how actively their cells move and rearrange.
@@ IV.2 Oscillatory Shear Reveals Cell Motility modulates Viscoelastic Crossover Frequency
To probe the viscoelastic properties of the tissue, we subjected the model tissue to low-amplitude oscillatory shear across a wide range of frequencies. This approach enabled us to extract the frequency-dependent storage modulus ($G^{\prime}$), representing the elastic (energy-storing) response, and the loss modulus ($G^{\prime\prime}$), representing the viscous (energy-dissipating) response. Together, these moduli provide a quantitative measure of the solid or fluid-like behavior of the tissue at different timescales of mechanical deformation.
At low oscillation frequencies, where the period of deformation exceeds the intrinsic stress relaxation time of the tissue, we observed that $G^{\prime\prime}>G^{\prime}$. This indicates a viscosity-dominated response, in which the tissue can fully relax the imposed stress before the next deformation cycle begins, making the tissue behave more like a fluid. This corresponds to situations where cells have sufficient time to have enough junctional rearrangements (T1 transitions), contributing to tissue flow (Fig. 2(d)).
In contrast, at high frequencies, the oscillation timescale becomes shorter than the tissue’s internal relaxation time, leading to an increase in elastic behavior of the tissue ($G^{\prime}>G^{\prime\prime}$, Fig. 2 (d)). Here, the tissue behaves more like a solid, storing deformation energy rather than dissipating it. The quick change of the shearing cycle does not give enough time for cellular junctional rearrangements or internal stress relaxation, causing the tissue to respond as an elastic material (Fig. 2(d)).
This fluid-to-solid transition is further quantified by the loss tangent $\tan\delta=\frac{G^{\prime\prime}}{G^{\prime}}$, which decreases with frequency [Fig.2(e)]. The characteristic crossover frequency, defined by the condition $G^{\prime\prime}=G^{\prime}$, marks the intrinsic relaxation timescale of the tissue. Importantly, we find that this crossover shifts to higher frequencies as cellular motility increases [Fig.2(f)], indicating that active cell movements facilitate more rapid stress relaxation.
This result highlights a key role for active cellular processes in modulating the viscoelastic landscape of the tissue, tuning its mechanical state between more fluid-like or solid-like regimes depending on the frequency of applied stress.
@@ V Temporally pulsated spatial perturbation generates permanent deformation in viscoelastic tissue
Having established the motility-dependent viscoelastic nature of the tissue, we next use the extracted viscoelastic timescales to predict tissue morphology under controlled external perturbations. Such analytical formulations offer a powerful means to uncover the fundamental principles underlying complex tissue dynamics. By abstracting the system into a continuum viscoelastic framework, we can extract scaling relations, isolate key parameter dependencies, and obtain predictions that remain broadly generalizable across contexts.
V.1 Analytical Treatment Shows Motility Driven Viscoelasticity and Perturbation Length Scale Controls Permanent Tissue Deformation
Motivated by the presence of multiple relaxation timescales in the tissue, we model the synthetic vertex tissue as a two-dimensional viscoelastic medium using a linear viscoelastic model with two timescalesKelly (2025). This representation comprises a dashpot with viscosity $\eta_{1}$ in series with a Kelvin–Voigt element, itself consisting of a spring with modulus $E$ in parallel with a dashpot of viscosity $\eta_{2}$ (Fig. 3(a) (I), (II) ). The constitutive stress–strain relation for this system is given by
@@
where $\tilde{\boldsymbol{\sigma}}$ and $\tilde{\boldsymbol{\varepsilon}}$ denote the stress and strain tensors, respectively. We then subject this material to an external perturbation that is both spatially patterned and temporally pulsed. Specifically, the forcing has a sinusoidal spatial profile of wavelength $k_{0}$ and is applied in periodic on–off cycles, with duration $\mathrm{T_{on}}$ and $\mathrm{T_{off}}$ respectively (Fig. 3(a) III):
@@
We found the deformation at the end of one cycle ($t=\mathrm{T_{on}}+\mathrm{T_{off}}$) is given in terms of the two timescales $\tau_{1}=\frac{\eta_{1}}{E}$ and $\tau_{2}=\frac{\eta_{2}}{E}$ as
@@
To directly connect the analytical deformation with experimentally accessible parameters such as motility, we first identify $\tau_{1}$ with the stress relaxation time, $\tau_{s}$. From Fig. 2(b), we established that $\tau_{s}$ scales inversely with motility, $\mathcal{M}$, i.e. $\tau_{s}\sim 1/\mathcal{M}$. Substituting this relation into the analytical expression yields a compact scaling form for the steady-state deformation amplitude:
@@
This expression reveals several key features. For any finite $\mathrm{T_{on}}$, the material exhibits a nonzero residual deformation at the end of each actuation cycle. Such residual deformations accumulate over successive cycles, potentially leading to lasting morphological transformations. In addition, the deformation amplitude scales linearly with the inverse of the fast relaxation timescale $\tau_{1}$. Finally, Relation S36 emphasizes the spatial dependence of the response: in the limit $k_{0}\to\infty$ (corresponding to short-wavelength perturbations), the deformation amplitude vanishes. Together, these results predict that the timescale of cellular relaxation and the wavelength of external forcing jointly determine the efficiency of morphogenetic remodeling.
@@ V.2 Active Vertex Model Tissue Validates Analytical Prediction
To test these analytical predictions, we applied to our synthetic tissue a perturbation similar to that described in Eq. 5. Each cycle consisted of an active phase of duration $\tau_{\text{on}}$ followed by a passive phase of duration $\tau_{\text{off}}$, chosen such that $\tau_{\text{on}}<\tau_{v}\ll\tau_{\alpha}$ and $\tau_{\text{off}}\lesssim\tau_{\mathrm{s}}$ (Fig. 3(b)). These parameters were selected as an extreme case relative to the analytical predictions to ensure that the simulated dynamics span the most separated timescales. This timing ensured that each active input was delivered within the tissue’s elastic response window, before complete stress relaxation. The forcing amplitude was kept below the intrinsic length scale of T1 transitions, thereby preventing the perturbation from directly inducing topological rearrangements. This setup allows us to isolate how short-lived external inputs applied in an elastic response window interact with slower internal relaxation processes to drive long-term tissue organization. Despite the short-lived nature of individual perturbations, our simulations show that tissues can undergo significant morphological changes over time, depending on both cell motility and the spatial scale of the perturbation.
@@
For long-wavelength perturbations ($\lambda\sim\frac{L_{y}}{4},L_{y}$ is the length of the tissue along y direction), tissue behavior shows strong dependence on motility. In low motility regimes ($\mathcal{M}\sim 0.00$), the tissue stress relaxation timescale ($\tau_{\mathrm{s}}$) is much longer than the relaxation period of the applied perturbation ($\tau_{\mathrm{off}}$) (Fig. 3(b)). Consequently, stress fails to dissipate neither via internal relaxation nor via junctional rearrangements. Instead, the asymmetric and cyclic nature of the sinusoidal forcing drives a coherent bulk rotation of the tissue, and no spatially structured patterns emerge. (Fig.  3(c)I). Intermediate motility ($\mathcal{M}=0.01$) allows some level of rearrangement but is insufficient to fully align with the imposed field, leading to partial morphological imprinting( Fig.  3(c) II). Only in high motility tissues ($\mathcal{M}=0.04$), where $\tau_{\mathrm{s}}\sim\mathrm{T_{off}}$ (Fig. 3(b)) junctional remodeling is efficient, and cells align locally with the imposed deformation(Fig.3(c)III) creating a strong morphogenetic pattern.
@@
When pulsatile perturbations are applied at short wavelengths ($\lambda\sim\tfrac{L_{y}}{16}$, roughly four cell diameters), the tissue fails to exhibit any long-term morphological adaptation across all motility levels. Unlike long-wavelength inputs, which impose global deformations, short-wavelength perturbations generate localized edge-length variations arising from the rapid spatial alternation of applied forces. These sharp local distortions increase the frequency of T1 transitions, allowing the tissue to rapidly relax back to its original configuration.
@@
In the low-motility regime, one might expect slow stress relaxation to favor shape retention, as in the long-wavelength case. However, here the imposed short-wavelength fluctuations directly drive local junctional rearrangements, overriding the slow viscoelastic response and preventing any stable morphological adaptation (Fig. 3(c)(IV)). For intermediate motility, junctional rearrangements occur even more rapidly, further accelerating local relaxation and again erasing any imposed pattern (Fig.3(c)(V)). At the highest motility levels, cells actively remodel their junctions, and the combination of intrinsic motility with edge-length fluctuations leads to transient small-scale boundary undulations. Yet, unlike the long-wavelength perturbations that generate coherent and lasting morphological changes, these fluctuations are quickly dissipated and do not translate into stable adaptations (Fig.3(c)(VI)).
@@
We further performed a quantitative comparison between analytical predictions and simulation results. Specifically, we computed the long-term deformation amplitude in Fourier space, $u_{n}(k_{0},\mathcal{M})$ (see S.I. Section S3 for details), and compared it against the theoretical scaling. The simulations show that the deformation amplitude increases linearly with motility $\mathcal{M}$, in agreement with the analytical prediction (Fig. 3(d)(I)). Also, both theory and simulations reveal a clear power-law dependence on perturbation length scale. While the theoretical prediction gives an exponent $\alpha=2$, the simulations yield a slightly smaller value, $\alpha\approx 1.34$ (Fig. 3(d)(II)), yet still capture the same underlying scaling behavior.

### v2

@@ IV Cell motility controls tissue viscoelasticity
Figure 2: Tissue rheology under standard mechanical protocols. (a) Stress relaxation response for tissues with different motility levels. The decay of shear stress $\sigma_{xy}(t)$ is shown following a step shear deformation. Higher motility leads to faster stress relaxation and lower residual stress. Inset: Long-time behavior of $\sigma_{xy}$ reveals a nonzero plateau, indicating residual stress retention. (b) Relaxation timescale $\tau_{\mathrm{s}}$ decreases with increasing motility, indicating enhanced fluidization. (c) Residual stress as a function of motility. More motile tissues retain less stress over time. (d) Frequency-dependent storage modulus $G^{\prime}$ and loss modulus $G^{\prime\prime}$. At low frequencies, viscous behavior dominates ($G^{\prime\prime}>G^{\prime}$); at higher frequencies, the tissue responds more elastically. (e) The loss tangent $\tan{\delta}=G^{\prime\prime}/G^{\prime}$ as a function of driving frequency $\omega_{0}$, illustrating the transition from viscous to elastic dominance. (f) Viscoelastic crossover time of the tissue($\tau_{V}$), defined by the crossover frequency where $G^{\prime}=G^{\prime\prime}$. $\tau_{V}$ shifts to lower values with increasing motility, reflecting faster stress relaxation dynamics in more active tissues.
@@
To further probe tissue rheology, we subjected the tissue to controlled shear deformations under varying levels of cellular motility. This approach allowed us to quantify fundamental viscoelastic properties — including stress-relaxation dynamics and the frequency-dependent storage and loss moduli — and thereby establish a direct link between local active fluctuations and the emergent tissue-scale mechanics.
@@
IV.1 Stress relaxation test reveals motility-driven residual stress retention
@@
The anterior–posterior variation in tissue viscosity suggests that different regions of the PSM may respond differently to mechanical loading. To investigate how this rheological heterogeneity affects stress dissipation, we performed stress-relaxation tests on tissues with varying motility levels. In each simulation, we applied a step shear by deforming the tissue affinely and fixing the outermost cell layers to maintain the applied strain. We then tracked the temporal evolution of shear stress in the bulk region, away from the boundary. By comparing tissues with low, medium, and high motility representing anterior, middle, and posterior PSM, we assessed how active cellular dynamics influence the tissue’s ability to dissipate stress.
@@
Across all motility regimes, we observed a characteristic biphasic relaxation response. At short to intermediate timescales, shear stress decayed rapidly [Fig. 2(a)], and the slope of this decay depended strongly on motility: higher motility produced steeper relaxation curves, indicating faster stress dissipation(Fig. 2(b)).
@@
At longer timescales, the stress curves plateaued, revealing a non-zero residual stress in the tissue [Fig. 2(a), Inset]. This behavior indicates solid-like features where cells cannot fully rearrange to relax all internal stresses. The magnitude of this residual stress was highest in the low-motility (anterior) tissues and decreased progressively with increasing motility (posterior) [Fig. 2(c)]. This motility-dependent decline in residual stress agrees qualitatively with experimental observations in embryonic PSM, where the posterior regions retain less residual stress than anterior ones kimEmbryonicTissuesActive2021a; mongeraFluidtosolidJammingTransition2018.
@@
Together, these stress-relaxation results suggest that motility not only governs the rate of stress dissipation but also tunes the mechanical state of the tissue from a more elastic, stress-retaining regime to a more fluidized, actively relaxing state. These findings help explain how developing tissues adjust their mechanical behavior during morphogenesis by tuning how actively their cells move and rearrange.
@@ IV.2 Oscillatory shear reveals cell motility modulates viscoelastic crossover frequency
To probe the viscoelastic properties of the tissue, we subjected the model tissue to low-amplitude oscillatory shear across a wide range of frequencies. This approach enabled us to extract the frequency-dependent storage modulus ($G^{\prime}$), representing the elastic (energy-storing) response, and the loss modulus ($G^{\prime\prime}$), representing the viscous (energy-dissipating) response. Together, these moduli quantify the relative solid- and fluid-like behavior of the tissue over different deformation timescales.
At low oscillation frequencies, where the period of deformation exceeds the intrinsic stress-relaxation time of the tissue, we observed that $G^{\prime\prime}>G^{\prime}$ (Fig. 2(d)). This indicates a viscosity-dominated response, in which the tissue can fully relax the imposed stress before the next deformation cycle begins, making the tissue behave more like a fluid. This corresponds to situations where cells have sufficient time to undergo junctional rearrangements (T1 transitions), enabling tissue flow (Fig. 2(d)).
In contrast, at high frequencies, the oscillation timescale becomes shorter than the tissue’s internal relaxation time, resulting in an elastic-dominated response with $G^{\prime}>G^{\prime\prime}$ [Fig. 2(d)]. Here, rapid shear cycle does not allow adequate time for cellular rearrangements or internal stress relaxation, causing the tissue to store rather than dissipate deformation energy.
This fluid-to-solid transition is further characterized by the loss tangent $\tan\delta=\frac{G^{\prime\prime}}{G^{\prime}}$, which decreases monotonically with frequency [Fig.2(e)]. The characteristic crossover frequency($\omega_{c}$), defined by the condition $G^{\prime\prime}(\omega_{c})=G^{\prime}(\omega_{c})$, marks the intrinsic relaxation timescale of the tissue. Interestingly, this crossover shifts to higher frequencies as cellular motility increases [Fig.2(f)], indicating that active cell motion accelerates stress relaxation and pushes the viscoelastic transition to faster timescales.
These results highlight the central role of active cellular processes in shaping the viscoelastic landscape of the tissue, tuning its mechanical state between fluid-like and solid-like regimes depending on both motility and deformation frequency.
@@ V Temporally pulsated spatial perturbation generates permanent deformation in viscoelastic tissue
Having established the motility-dependent viscoelastic nature of the tissue, we next use the extracted viscoelastic timescales to predict tissue morphology under controlled external perturbations analytically. Such analytical formulations offer a powerful means to uncover the fundamental principles governing complex tissue dynamics. By abstracting the system into a continuum viscoelastic framework, we can extract scaling relations, isolate key parameter dependencies, and obtain predictions that remain broadly generalizable across contexts.
V.1 Analytical treatment shows motility-driven viscoelasticity and perturbation length scale controls permanent tissue deformation
Motivated by the presence of multiple relaxation timescales in the tissue, we model the synthetic vertex tissue as a two-dimensional viscoelastic material using a linear viscoelastic model with two distinct relaxation timescales kellyMechanicsLectureNotes2025. This representation consists of a dashpot with viscosity $\eta_{1}$ in series with a Kelvin–Voigt element comprising a spring of modulus $E$ in parallel with a dashpot of viscosity $\eta_{2}$ [Fig. 3(a) (I), (II)]. The resulting constitutive relation between stress and strain is
@@
where $\tilde{\boldsymbol{\sigma}}$ and $\tilde{\boldsymbol{\varepsilon}}$ denote the stress and strain tensors, respectively. We then subject this material to an external perturbation that is both spatially patterned and temporally pulsed. Specifically, the forcing has a sinusoidal spatial profile of wavelength $k_{0}$ and is applied in periodic on-off cycles of durations $\mathrm{T_{on}}$ and $\mathrm{T_{off}}$:
@@
as illustrated in Fig. 3(a)(III). Solving the constitutive equation under this forcing yields the deformation at the end of a single on–off cycle ($t=\mathrm{T_{on}}+\mathrm{T_{off}}$), expressed in terms of the two characteristic relaxation times $\tau_{1}=\eta_{1}/E$ and $\tau_{2}=\eta_{2}/E$:
@@
To connect this analytical deformation with experimentally accessible parameters such as motility, we identify $\tau_{1}$ with the stress-relaxation time, $\tau_{s}$. From Fig. 2(b), we established that $\tau_{s}$ scales inversely with motility, $\mathcal{M}$, i.e. $\tau_{s}\sim 1/\mathcal{M}$. Substituting this relation into Eq. 7 yields a compact scaling form for the steady-state deformation amplitude:
@@
This scaling relation highlights several key features. First, for any finite $\mathrm{T_{on}}$, the material retains a nonzero residual deformation after each actuation cycle, enabling cumulative and potentially permanent shape changes. Second, the deformation amplitude increases linearly with the inverse of the fast relaxation timescale $\tau_{1}$, indicating that faster-relaxing tissues respond more strongly to pulsatile forcing. Finally, Eq. S36 emphasizes the strong spatial dependence of the response: in the limit $k_{0}\to\infty$ (short-wavelength perturbations), the deformation amplitude vanishes. Together, these results demonstrate that the interplay between cellular relaxation timescales and the spatial wavelength of external forcing jointly determines the efficiency of morphogenetic remodeling.
@@ V.2 Active vertex model tissue validates analytical prediction
To test these analytical predictions, we subjected the synthetic tissue to a spatially patterned pulsed forcing identical in form to Eq. 5. Each cycle consisted of an active phase of duration $\mathrm{T_{on}}$ followed by a passive phase of duration $\mathrm{T_{off}}$, chosen such that $\mathrm{T_{on}}<\tau_{V}\ll\tau_{\alpha}$ and $\mathrm{T_{off}}\lesssim\tau_{\mathrm{s}}$ [Fig. 3(b)]. These parameters were selected as an extreme regime relative to the analytical predictions to ensure that the simulated dynamics span well-separated timescales. This choice ensures that each active input pulse is delivered within the tissue’s elastic response window, before complete stress relaxation. The forcing amplitude was kept below the intrinsic length scale of T1 transitions, thereby preventing the perturbation from directly inducing topological rearrangements. Consequently, the setup isolates how rapid, transient inputs interact with slower internal viscoelastic relaxation to generate long-term morphological outcomes. Despite the short-lived nature of each pulse, our simulations show that tissues can undergo significant morphological changes over time, depending on both cell motility and the spatial scale of the perturbation.
@@
For long-wavelength perturbations ($\lambda\sim\frac{L_{y}}{4},L_{y}$ is the length of the tissue along the y direction), tissue response exhibits a pronounced dependence on motility. In low motility regimes ($\mathcal{M}\sim 0.00$), the tissue stress relaxation timescale ($\tau_{\mathrm{s}}$) is much longer than the relaxation period of the applied perturbation ($\mathrm{T_{off}}$) (Fig. 3(b)). Consequently, stress fails to dissipate either via internal relaxation or via junctional rearrangements. Instead, the asymmetric and cyclic nature of the sinusoidal forcing drives a coherent bulk rotation of the tissue (a symmetric forcing eliminates bulk rotation), and no spatially structured patterns emerge [Fig. 3(c)I]. At intermediate motility ($\mathcal{M}=0.01$), isolated rearrangements occur, enabling partial imprinting of the pattern [Fig. 3(c)II]. Only at high motility ($\mathcal{M}=0.04$), where $\tau_{\mathrm{s}}\sim\mathrm{T_{off}}$ (Fig. 3(b)), tissue efficiently remodel via T1 transitions. In this regime, cells locally align with the imposed deformation, producing a clear and persistent morphogenetic pattern [Fig. 3(c)III].
@@
By contrast, short-wavelength perturbations ($\lambda\sim\tfrac{L_{y}}{16}$, roughly four cell diameters) fail to generate long-term morphological adaptation across all motility values. Unlike long-wavelength forcing, which imposes global deformations, short-wavelength perturbations generate localized edge-length variations arising from the rapid spatial alternation of applied forces. These sharp localized distortions trigger frequent T1 transitions, allowing the tissue to relax back to its initial state on timescales much shorter than those required for global deformation.
@@
In the low-motility regime, one might expect slow stress relaxation to favor shape retention, as in the long-wavelength case. However, here the imposed short-wavelength fluctuations directly drive local junctional rearrangements, overriding the slow viscoelastic response and preventing any stable morphological adaptation [Fig. 3(c)(IV)]. For intermediate motility, junctional rearrangements occur even more rapidly, further accelerating local relaxation and again erasing any imposed pattern [Fig.3(c)(V)]. At the highest motility levels, cells actively remodel their junctions, and the combination of intrinsic motility with edge-length fluctuations leads to transient small-scale boundary undulations. Yet, unlike the long-wavelength perturbations that generate coherent and lasting morphological changes, these fluctuations are quickly dissipated and do not translate into stable adaptations [Fig.3(c)(VI)].
@@
Finally, we performed a quantitative comparison between analytical predictions and simulations by measuring the long-term deformation amplitude in Fourier space, $u_{n}(k_{0},\mathcal{M})$ (see S.I. Sec. S3). The simulations show a linear increase in deformation amplitude with motility $\mathcal{M}$, in agreement with the analytical scaling (Fig. 3(d)I). Both theory and simulations reveal a clear power-law dependence on perturbation length scale. While the theoretical prediction gives an exponent $\alpha=2$, the simulations yield a slightly smaller value, $\alpha\approx 1.34$ [Fig. 3(d)(II)]. Despite this quantitative difference, the underlying scaling behavior is preserved, validating the key analytical predictions.